Hausmitteilung

An alle Mitarbeiterinnen und Mitarbeiter Institut für Kern- und Hadronenphysik

Name

Telefon (03 51) 2 60-99 99 Telefax (03 51) 2 60-37 00 V.Name@fz-rossendorf.de

16. Januar 2001

Sehr geehrte ...

Text

signname